# The frame-dragging vector potential on galaxy scales from DM-only Newtonian $N$-body simulations

### v2

@@ 3 The Illustris-TNG simulations
As already mentioned in the introduction, the behaviour of the frame-dragging potential is poorly known at very small scales, where non-linearities are expected to generate rotational momentum flows of matter. Although barrera-hinojosaVectorModesLCDM2021 probed the vector potential inside dark matter halos with masses $10^{12.5}\text{--}10^{15}\;h^{-1}\,\mathrm{M}_{\odot}$, thanks to the great resolution capabilities of the AMR technique, they focused on a large cosmological simulation of sole dark matter with a box size of $256\;h^{-1}$ Mpc and a particle mass resolution of $1.33\times 10^{9}\;h^{-1}\,\mathrm{M}_{\odot}$ (the same mass resolution was used by adamekGeneralRelativityCosmic2016, for larger simulations). On the other hand, thomasFullyNonlinearPostFriedmann2015 analysed smaller simulations ranging from $80$ to $320\;h^{-1}$ Mpc, but with mass resolution greater than $3.97\times 10^{8}\;h^{-1}\,\mathrm{M}_{\odot}$ in all cases.
@@
On the other hand, the dark-matter-only versions of these simulations are particularly useful for studying the non-linear evolution of dark matter halos and their substructures, gravitational potentials, and cosmic velocity fields, without the complicating effects of baryonic physics. Table (1) lists the main parameters of the simulations analysed in this paper.
@@
Table 1: Details of the IllustrisTNG simulations used in this paper: the box size, the number of dark matter particles $N_{\mathrm{DM}}$, and the mass of each dark matter particle $m_{\mathrm{DM}}$.
@@
In the smallest simulation, i.e. TNG50-2-Dark, dark matter particles have a mass of $2.9\times 10^{6}\;h^{-1}\,\mathrm{M}_{\odot}$, meaning that the mass resolution is at least two orders of magnitude better than the studies mentioned above (bruniComputingGeneralrelativisticEffects2014a; thomasFullyNonlinearPostFriedmann2015; thomasGravityNonlinearScales2015; barrera-hinojosaVectorModesLCDM2021; adamekGeneralRelativityCosmic2016). On the other hand, the largest simulation is used to ensure consistency with these works. Higher-resolution simulations, such as TNG50-1-Dark, provide finer details but will not be employed due to the prohibitive computational cost of extracting the relevant fields.
The simulations have been evolved from redshift $z=127$ and assuming the Planck20152015 cosmology (planckcollaborationPlanck2015Results2016), with a Hubble parameter $h=0.6774$, matter density parameter $\Omega_{\mathrm{m}}=0.3089$, and dark energy density parameter $\Omega_{\Lambda}=0.6911$. Multiple snapshots from $z=20$ to $z=0$ are available, in order to investigate the time evolution of the vector potential. In future works, the use of the IllustrisTNG suite will also straightforwardly facilitate the extension to the hydrodynamical counterpart of the simulations in order to investigate the role of baryons in this context, which has yet to be evaluated in the literature.
@@ 4 Extracting the gravito-magnetic potential
A common approach in cosmological simulations is grid-based interpolation, such as the Cloud-in-Cell (CIC) and Triangular-Shaped Cloud (TSC) methods (Hockney-Eastwood:1988). These techniques work by assigning particle properties to a fixed spatial grid and averaging over nearby points to generate a continuous field representation. While efficient and straightforward to implement, such methods fail to accurately capture small-scale structures, as they can lead to over-smoothing.
@@ 4.1 The Delaunay Tessellation Field Estimator
A more adaptive and accurate alternative is the Delaunay Tessellation Field Estimator (DTFE, schaapContinuousFieldsDiscrete2000), which reconstructs continuous fields using a Delaunay triangulation of the particle positions. This technique partitions the simulation domain into a unique tessellation of tetrahedra, whose vertices are defined by the particle positions and that do not contain any particle, and then draws the field on a regular grid. In detail, the density field is first evaluated on each vertex as inversely proportional to the volume of the contiguous tetrahedra, while the velocity field is already known. Then, for each tetrahedron, the field is linearly interpolated on $n$ random sampling points extracted inside the tetrahedron. [^3] In this paper the standard value of $n=100$ is adopted. Finally, the field at each grid cell is computed as the volume-weighted average of the field over the entire volume of the grid cell, where the contribution of each sample point is distributed to the grid cell in which it resides.
@@
The DTFE is a widely used method in cosmological simulations for extracting continuous fields and their derivatives from discrete particle data. Unlike fixed-grid methods, it dynamically adapts to the local density of particles, providing higher resolution in dense regions such as halos and filaments, while maintaining coarser resolution in cosmic voids. This reduces artificial smoothing effects, such as assigning artificial zero velocities in voids, which is the major limitation of methods like CIC (pueblasGenerationVorticityVelocity2009). Moreover, the volume averaging of the fields reduces the Poisson noise. This adaptability makes the DTFE particularly suited for estimating both the density and velocity fields in highly non-linear environments. The DTFE also enables direct computation of spatial derivatives, including velocity divergence and vorticity, without requiring additional smoothing or convolution steps.
@@
However, the method is not without limitations. The computational cost of constructing the tessellation increases significantly with the number of particles, whereas the lack of a straightforward deconvolution of the window function prevents the correction for interpolation effects at very small scales. Additionally, the method’s reliability can be affected in regions with caustics or orbit-crossing events, where discontinuities in the underlying fields may lead to inaccuracies in the computed gradients (hahnPropertiesCosmicVelocity2015). Here, in fact, the velocity divergence and vorticity estimated via the finite difference scheme of DTFE are not completely reliable, whereas phase-space interpolation methods are expected to perform better (Abel2012MNRAS.427...61A).
@@
In this paper, the publicly available DTFE implementation by cautunDTFEPublicSoftware2019a is employed to extract the density and velocity fields and their gradients. All the fields are interpolated into a cubic grid with $N_{\mathrm{grid}}=1024^{3}$ nodes. This number ensures a good convergence of the results, as detailed in Appendix A.1.
@@
In the following, in order to compare with $\delta$ and the velocity $\mathbf{v}$, we find convenient to work with the “momentum field" $\mathbf{p}$ defined as thomasFullyNonlinearPostFriedmann2015
@@ 4.2 Fourier analysis
The spatial properties of the extracted fields are then analysed in Fourier space, where power spectra provide a compact statistical description of how structures are distributed across scales. Following the convention below, the power spectrum of a vector field $\mathbf{v}$ is defined as the sum of the power spectra of its components:
@@
The computation of the power spectra is performed through the python library nbodykit (handNbodykitOpensourceMassively2018), where the Fourier coefficients are averaged over spherical shells in $k$-space. The width of the bins is set to $4\pi/L$, with $L$ denoting the size of the simulation box. All power spectra are truncated at the Nyquist wave number, namely $k_{\mathrm{max}}=\pi N_{\mathrm{grid}}^{1/3}/L$, while the minimum wave number is set to the fundamental mode $k_{\mathrm{box}}=2\pi/L$. For the gravitational potentials, the results are expressed in terms of the dimensionless power spectra, defined as
@@
This procedure is more robust than computing the divergence and curl in real space and then Fourier transforming them, particularly in the presence of discontinuities or caustics (hahnPropertiesCosmicVelocity2015).
@@
To interpret the amplitude and scale dependence of the vector potential, it is useful to compare these power spectra with analytical expectations from perturbation theory. Here, the vector potential appears as a second-order term as it is sourced by the product of first-order density and velocity divergence, in the case of the CDM fluid. The power spectrum of the vector potential at the second-order approximation is given by luCosmologicalBackgroundVector2009, which is equivalent to the integral in the form provided by barrera-hinojosaVectorModesLCDM2021:
@@
The velocity potential $\mathit{v}$ is defined so that its Laplacian is equal to the velocity divergence, i.e. $\nabla^{2}\mathit{v}=\nabla\cdot\mathbf{v}$. In this way, according to the continuity equation (27) in linear theory, the power spectrum of the velocity potential is determined by the matter power spectrum, namely $\Delta_{\mathit{v}}=k^{-4}\Delta_{\delta}$. The reader must note that Equation (17) differs from Equation 30 in barrera-hinojosaVectorModesLCDM2021 by a factor $a^{2}$. This is due to the different definition of the vector potential $\mathbf{B}$, already mentioned in Section 2.1.
@@ 4.2.1 Global solution
Alternatively, Equations (9) and (6) can be globally solved in the Fourier space. For the scalar potential, the Poisson equation in Fourier space becomes
@@
These equations allow the derivation of $\phi$ and $\mathbf{B}$ globally across the entire simulation volume, taking advantage of periodic boundary conditions. These are defined up to an adding constant, therefore the two potentials are conventionally normalized in the real space so that they have zero mean in the simulation box, i.e. $\langle\phi\rangle=0$ and $\langle B_{i}\rangle=0$. This is consistent with the simulation box being a sample of an isotropic FLRW universe. The Fourier transforms are numerically computed through the FFT algorithm implemented in numpy (harris2020numpy), where the periodic boundary conditions are automatically incorporated. However, the inverse Fourier transform introduces spurious quantities at the caustic locations, where the velocity field is discontinuous and the derivative is infinite (hahnPropertiesCosmicVelocity2015). Therefore, one should not expect this method to give accurate quantitative information, but at least a qualitative picture of the regions with high gravito-magnetic potential.
@@ 5.1.1 Matter density
This is a well-known limitation of cosmological $N$-body simulations performed on such small volumes. For instance, Bagla2009 suggests that the size of the simulation volume should be chosen so that the amplitude of fluctuations at the box scale (and at larger scales) is ignorable, or equivalently the starting redshift (Power2003), while adamekGeneralRelativityCosmic2016 found that in order to get a good agreement between simulation results and perturbation theory, the box should be large enough to contain the matter–radiation equality scale, which is $k_{\mathrm{eq}}\approx 0.010\;h^{-1}$ Mpc (Planck2019). In the next sections, we will see how this problem affects the calculation of the gravito-magnetic potential on these scales even more clearly and how to deal with it correctly.
At small scales, the matter power spectra deviate from the HaloFit prediction due to the limited spatial resolution of the DTFE and its window function. It is important to ensure that our numerical results are robust and independent of the simulation parameters, like the mass resolution, the box size, as well as the grid resolution of the tessellation algorithm. Such convergence checks are presented in Appendix A, where a DTFE grid of $1024^{3}$ points is found to provide good convergence for the particle-mesh interpolation. Although not shown here, it has been confirmed that the Cloud-in-Cell assignment performs much worse than DTFE for the same grid resolution, as expected (pueblasGenerationVorticityVelocity2009; thomasFullyNonlinearPostFriedmann2015).
@@ 5.1.2 Velocity and momentum density
At large scales, the power spectrum of the momentum field $\mathbf{p}$ overlaps the power spectrum of the velocity field (Figure 2). This is expected in linear theory since the term $\delta\mathbf{v}$ behaves as a second-order perturbation, while at smaller scales, this term becomes relevant as a result of the non-linear evolution.
@@ 5.1.3 Vorticity and divergence
As expected, the velocity divergence shows good agreement with linear theory at large scales, while exhibiting a well-known dip at around $k\approx 1.5\;h$ Mpc-1 for the 205 and 75 $h^{-1}$ Mpc simulations, and at around $k\approx 3\;h$ Mpc-1 for the 35 $h^{-1}$ Mpc simulation (barrera-hinojosaVectorModesLCDM2021).
@@
For the velocity vorticity, the power spectrum exhibits a peak. The peak is located at around the same scale the divergence shows the dip, i.e. $k\approx 1.2\;h$ Mpc-1, consistently with barrera-hinojosaVectorModesLCDM2021 and jelic-cizmekGenerationVorticityCosmological2018, with the smaller simulation showing a peak slightly shifted towards smaller scales, ie $k\approx 2.5\;h$ Mpc-1.
@@
This suggests that as non-linear structures form, a fraction of the power in the velocity divergence is transferred to vorticity, meaning that the conservation of angular momentum prevents further infall of matter and forces the particles to rotate around structures (jelic-cizmekGenerationVorticityCosmological2018).
In the context of cold dark matter, the pressure-less perfect fluid is described in the single-stream regime. In this approximation, at each point in space, the velocity field is uniquely defined, that is to say dark matter particles move coherently without crossing each other’s trajectories. Under this assumption, vorticity is expected to vanish at all orders of perturbation theory (luCosmologicalBackgroundVector2009).
@@
However, in $N$-body simulations, as structures grow and gravitational collapse becomes non-linear, the single-stream approximation breaks down due to orbit crossing. In this multi-stream regime, particles with different velocities can coexist at the same spatial location, leading to the generation of vorticity (pueblasGenerationVorticityVelocity2009).
The emergence of vorticity in simulations is strongly dependent on the mass resolution (jelic-cizmekGenerationVorticityCosmological2018). If the resolution is too low, the fine details of orbit crossing may not be well captured, leading to an underestimation of vorticity. Conversely, a higher resolution allows for a more accurate representation of the small-scale velocity field, and a better resolution of the transition to the multi-stream regime. In Appendix A.2, three different mass resolutions for the simulation TNG50-Dark are tested. Although, at small scales, a coarser resolution attributes more spurious power, the shape and amplitude of the vorticity power spectrum are preserved. For the momentum vorticity, such dependence on the mass resolution is weaker, as well as for the divergence of both fields.
@@
The power spectra of the last two terms are plotted in Figure 4. In agreement with several previous works (bruniComputingGeneralrelativisticEffects2014a; thomasFullyNonlinearPostFriedmann2015; barrera-hinojosaVectorModesLCDM2021), the main component that sources the momentum vorticity is $\nabla\delta\times\mathbf{v}$, while the velocity vorticity alone is subdominant, as expected being $\delta\gg 1$.

### v3

@@ 3 The Illustris-TNG simulations
As already mentioned in the introduction, the behaviour of the frame-dragging potential is poorly known at very small scales, where non-linearities are expected to generate rotational momentum flows of matter. Although Barrera-Hinojosa et al. (2021) probed the vector potential inside dark matter halos with masses $10^{12.5}\text{--}10^{15}\;h^{-1}\,\mathrm{M}_{\odot}$, thanks to the great resolution capabilities of the AMR technique, they focused on a large cosmological simulation of sole dark matter with a box size of $256\;h^{-1}$ Mpc and a particle mass resolution of $1.33\times 10^{9}\;h^{-1}\,\mathrm{M}_{\odot}$ (the same mass resolution was used by Adamek et al., 2016, for larger simulations). On the other hand, Thomas et al. (2015b) analysed smaller simulations ranging from $80$ to $320\;h^{-1}$ Mpc, but with mass resolution greater than $3.97\times 10^{8}\;h^{-1}\,\mathrm{M}_{\odot}$ in all cases.
@@
On the other hand, the dark-matter-only versions of these simulations are particularly useful for studying the non-linear evolution of dark matter halos and their substructures, gravitational potentials, and cosmic velocity fields, without the complicating effects of baryonic physics. In this case, the dark matter particles are more massive than in the case with baryons, precisely to compensate for the lack of baryons and preserve the matter density of the universe. Table (1) lists the main parameters of the dark-matter-only simulations analysed in this paper.
@@
Table 1: Details of the IllustrisTNG simulations used in this paper: the box size, the number of dark matter particles $N_{\mathrm{DM}}$, and the mass of each dark matter particle $m_{\mathrm{DM}}$. Simulations used for the main results are highlighted in bold; those in plain text were only used for the convergence tests in Appendix.
@@
In the smallest simulation, i.e. TNG50-2-Dark, dark matter particles have a mass of $2.9\times 10^{6}\;h^{-1}\,\mathrm{M}_{\odot}$, meaning that the mass resolution is at least two orders of magnitude better than the studies mentioned above (Bruni et al., 2014; Thomas et al., 2015b, a; Barrera-Hinojosa et al., 2021; Adamek et al., 2016). On the other hand, the largest simulation is used to ensure consistency with these works. Higher-resolution simulations, such as TNG50-1-Dark, provide finer details but will not be employed due to the prohibitive computational cost of extracting the relevant fields.
The IllustrisTNG simulations were evolved from redshift $z=127$ and assuming the Planck20152015 cosmology (Planck Collaboration et al., 2016) [^3] Note that the Planck2018 cosmology was not yet available when the IllustrisTNG project started., with a Hubble parameter $h=0.6774$, matter density parameter $\Omega_{\mathrm{m}}=0.3089$, and dark energy density parameter $\Omega_{\Lambda}=0.6911$. Multiple snapshots from $z=20$ to $z=0$ are available on request, and we used them in order to investigate the time evolution of the vector potential. In future works, the use of the IllustrisTNG suite will also straightforwardly facilitate the extension to the hydrodynamical counterpart of the simulations in order to investigate the role of baryons in this context, which has yet to be evaluated in the literature.
@@ 4 Extracting the gravito-magnetic potential
A common approach in cosmological simulations is grid-based interpolation, such as the Cloud-in-Cell (CIC) and Triangular-Shaped Cloud (TSC) methods (Hockney and Eastwood, Inc., Bristol, PA, USA, 1988). These techniques work by assigning particle properties to a fixed spatial grid and averaging over nearby points to generate a continuous field representation. While efficient and straightforward to implement, such methods fail to accurately capture small-scale structures, as they can lead to over-smoothing.
@@ 4.1 The Delaunay Tessellation Field Estimator
As already implemented by Bruni et al. (2014) and Thomas et al. (2015b), a more adaptive and accurate alternative is the Delaunay Tessellation Field Estimator (Schaap and van de Weygaert, 2000), which reconstructs continuous fields using a Delaunay triangulation of the particle positions. This technique partitions the simulation domain into a unique tessellation of tetrahedra, whose vertices are defined by the particle positions and that do not contain any particle, and then draws the field on a regular grid. In detail, the density field is first evaluated on each vertex as inversely proportional to the volume of the contiguous tetrahedra, while the velocity field is already known since the particle velocity is given as input. Then, for each tetrahedron, the field is linearly interpolated on $n$ random sampling points extracted inside the tetrahedron. [^4] In this paper the standard value of $n=100$ is adopted. We did not check the results for different values, as Bruni et al. (2014) varied $n$ up to 1000 and found no difference to the results. Finally, the field at each grid cell is computed as the volume-weighted average of the field over the entire volume of the grid cell, where the contribution of each sample point is distributed to the grid cell in which it resides.
@@
The DTFE is a widely used method in cosmological simulations for extracting continuous fields and their derivatives from discrete particle data. Unlike fixed-grid methods, it dynamically adapts to the local density of particles, providing higher resolution in dense regions such as halos and filaments, while maintaining coarser resolution in cosmic voids. This reduces artificial smoothing effects, such as assigning artificial zero velocities in voids, which is the major limitation of methods like CIC (Pueblas and Scoccimarro, 2009). Moreover, the volume averaging of the fields reduces the Poisson noise. This adaptability makes the DTFE particularly suited for estimating both the density and velocity fields in highly non-linear environments. The DTFE also enables direct computation of spatial derivatives, including velocity divergence and vorticity, without requiring additional smoothing or convolution steps.
@@
However, the method is not without limitations. The computational cost of constructing the tessellation increases significantly with the number of particles, whereas the lack of a straightforward deconvolution of the window function prevents the correction for interpolation effects at very small scales. Additionally, the method’s reliability can be affected in regions with caustics or orbit-crossing events, where discontinuities in the underlying fields may lead to inaccuracies in the computed gradients (Hahn et al., 2015). Here, in fact, the velocity divergence and vorticity estimated via the finite difference scheme of DTFE are not completely reliable, whereas phase-space interpolation methods are expected to perform better (Abel2012MNRAS.427...61A).
@@
In this paper, the publicly available DTFE implementation by Cautun and van de Weygaert (2019) is employed to extract the density and velocity fields and their gradients, following the same methodology of Bruni et al. (2014) and Thomas et al. (2015b). All the fields are interpolated into a cubic grid with $N_{\mathrm{grid}}=1024^{3}$ nodes. This number ensures a good convergence of the results, as detailed in Appendix A.1.
@@
In the following, in order to compare with $\delta$ and the velocity $\mathbf{v}$, we find convenient to work with the “momentum field" $\mathbf{p}$ defined as Thomas et al. (2015b)
@@ 4.2 Fourier analysis
In order to solve Equation (9), Bruni et al. (2014) and Thomas et al. (2015b) proposed to derive the properties of the extracted fields in Fourier space, where power spectra provide a compact statistical description of how structures are distributed across scales. Following the convention below, the power spectrum of a vector field $\mathbf{v}$ is defined as the sum of the power spectra of its components:
@@
The computation of the power spectra is performed through the python library nbodykit (Hand et al., 2018), where the Fourier coefficients are averaged over spherical shells in $k$-space. The width of the bins is set to $4\pi/L$, with $L$ denoting the size of the simulation box. All power spectra are truncated at the Nyquist wave number, namely $k_{\mathrm{max}}=\pi N_{\mathrm{grid}}^{1/3}/L$, while the minimum wave number is set to the fundamental mode $k_{\mathrm{box}}=2\pi/L$. For the gravitational potentials, the results are expressed in terms of the dimensionless power spectra, defined as
@@
This procedure is more robust than computing the divergence and curl in real space and then Fourier transforming them, particularly in the presence of discontinuities or caustics (Hahn et al., 2015).
@@
To interpret the amplitude and scale dependence of the vector potential, it is useful to compare these power spectra with analytical expectations from perturbation theory. Here, the vector potential appears as a second-order term as it is sourced by the product of first-order density and velocity divergence, in the case of the CDM fluid. The power spectrum of the vector potential at the second-order approximation is given by Lu et al. (2009), which is equivalent to the integral in the form provided by Barrera-Hinojosa et al. (2021, Equation 30):
@@
The velocity potential $\mathit{v}$ is defined so that its Laplacian is equal to the velocity divergence, i.e. $\nabla^{2}\mathit{v}=\nabla\cdot\mathbf{v}$. In this way, according to the continuity equation (27) in linear theory, the power spectrum of the velocity potential is determined by the matter power spectrum, namely $\Delta_{\mathit{v}}=k^{-4}\Delta_{\delta}$. The reader must note that Equation (17) differs from Equation 30 in Barrera-Hinojosa et al. (2021) by a factor $a^{2}$. This is due to the different definition of the vector potential $\mathbf{B}$, already mentioned in Section 2.1.
@@ 4.2.1 Global solution
Alternatively to the power spectrum method presented by Bruni et al. (2014) and Thomas et al. (2015b), Equations (9) and (6) can also be globally solved in the Fourier space. Therefore, contrary to previous analyses, this allows us to reconstruct the solution over the entire volume of the simulation. For the scalar potential, the Poisson equation in Fourier space becomes
@@
These equations allow the derivation of $\phi$ and $\mathbf{B}$ globally across the entire simulation volume, taking advantage of periodic boundary conditions. These are defined up to an adding constant, therefore the two potentials are conventionally normalized in the real space so that they have zero mean in the simulation box, i.e. $\langle\phi\rangle=0$ and $\langle B_{i}\rangle=0$. This is consistent with the simulation box being a sample of an isotropic FLRW universe. The Fourier transforms are numerically computed through the FFT algorithm implemented in numpy (harris2020numpy), where the periodic boundary conditions are automatically incorporated. However, the inverse Fourier transform introduces spurious quantities at the caustic locations, where the velocity field is discontinuous and the derivative is infinite (see Hahn et al., 2015, for an extended comparison between spectral and real space derivatives). Since such spurious quantities are typically localized at the edges of filaments, these are smoothed out by the $k^{2}$ factor of the elliptic equation (22). Therefore, even though one should not expect this method to give accurate quantitative information near caustics, we expect a minimal impact on the overall solution of the gravito-magnetic potential.
@@ 5.1.1 Matter density
This is a well-known limitation of cosmological $N$-body simulations performed on such small volumes. For instance, Bagla2009 suggests that the size of the simulation volume should be chosen so that the amplitude of fluctuations at the box scale (and at larger scales) is ignorable, or equivalently the starting redshift (Power2003), while Adamek et al. (2016) found that in order to get a good agreement between simulation results and perturbation theory, the box should be large enough to contain the matter–radiation equality scale, which is $k_{\mathrm{eq}}\approx 0.010\;h^{-1}$ Mpc (Planck2019). In the next sections, we will see how this problem affects the calculation of the gravito-magnetic potential on these scales even more clearly and how to deal with it correctly.
At small scales, the matter power spectra deviate from the HaloFit prediction due to the limited spatial resolution of the DTFE and its window function. It is important to ensure that our numerical results are robust and independent of the simulation parameters, like the mass resolution, the box size, as well as the grid resolution of the tessellation algorithm. Such convergence checks are presented in Appendix A, where a DTFE grid of $1024^{3}$ points is found to provide good convergence for the particle-mesh interpolation. Although not shown here, it has been confirmed that the Cloud-in-Cell assignment performs much worse than DTFE for the same grid resolution, as expected (Pueblas and Scoccimarro, 2009; Thomas et al., 2015b).
@@ 5.1.2 Velocity and momentum density
At large scales, the power spectrum of the momentum field $\mathbf{p}$ matches the power spectrum of the velocity field (Figure 2). This is expected in linear theory since the term $\delta\mathbf{v}$ behaves as a second-order perturbation, while at smaller scales, this term becomes relevant as a result of the non-linear evolution.
@@ 5.1.3 Vorticity and divergence
As expected, the velocity divergence shows good agreement with linear theory at large scales, while exhibiting a dip at around $k\approx 1.5\;h$ Mpc-1 for the 205 and 75 $h^{-1}$ Mpc simulations, and at around $k\approx 3\;h$ Mpc-1 for the 35 $h^{-1}$ Mpc simulation (Barrera-Hinojosa et al., 2021).
@@
For the velocity vorticity, the power spectrum exhibits a peak. The peak is located at around the same scale the divergence shows the dip, i.e. $k\approx 1.2\;h$ Mpc-1, consistently with Barrera-Hinojosa et al. (2021) and Jelic-Cizmek et al. (2018), with the smaller simulation showing a peak slightly shifted towards smaller scales, ie $k\approx 2.5\;h$ Mpc-1. The dip and the peak in the velocity divergence and vorticity power spectra, respectively, being at the similar position, have been interpreted as the consequence of shell crossing occurring around that scales, where the angular momentum can be large enough to inhibit structure growth as particles are forced into orbits rather than falling inward (Barrera-Hinojosa and Li, 2020; Jelic-Cizmek et al., 2018).
@@
This suggests that as non-linear structures form, a fraction of the power in the velocity divergence is transferred to vorticity, meaning that the conservation of angular momentum prevents further infall of matter and forces the particles to rotate around structures (Jelic-Cizmek et al., 2018).
In the context of cold dark matter, the pressure-less perfect fluid is described in the single-stream regime. In this approximation, at each point in space, the velocity field is uniquely defined, that is to say dark matter particles move coherently without crossing each other’s trajectories. Under this assumption, vorticity is expected to vanish at all orders of perturbation theory (Lu et al., 2009).
@@
However, in $N$-body simulations, as structures grow and gravitational collapse becomes non-linear, the single-stream approximation breaks down due to orbit crossing. In this multi-stream regime, particles with different velocities can coexist at the same spatial location, leading to the generation of vorticity (Pueblas and Scoccimarro, 2009).
The emergence of vorticity in simulations is strongly dependent on the mass resolution (Jelic-Cizmek et al., 2018). If the resolution is too low, the fine details of orbit crossing may not be well captured, leading to an underestimation of vorticity. Conversely, a higher resolution allows for a more accurate representation of the small-scale velocity field, and a better resolution of the transition to the multi-stream regime. In Appendix A.2, three different mass resolutions for the simulation TNG50-Dark are tested. Although, at small scales, a coarser resolution attributes more spurious power, the shape and amplitude of the vorticity power spectrum are preserved. For the momentum vorticity, such dependence on the mass resolution is weaker, as well as for the divergence of both fields.
@@
The power spectra of the last two terms are plotted in Figure 4. In agreement with several previous works (Bruni et al., 2014; Thomas et al., 2015b; Barrera-Hinojosa et al., 2021), the main component that sources the momentum vorticity is $\nabla\delta\times\mathbf{v}$, while the velocity vorticity alone is subdominant, as expected being $\delta\gg 1$.